What am I doing wrong here?

Statistical Methods 1 Online $19 Wade leanette Willharns & 1/23 Remaining: 00 293 This Test est: Chapters 1-3 Test Tme is Question: 1 pt 14 of 14 (14 complete) show Work contains the area of all the New Jersey counties in square mles 130 569 234 47 192 19 423 365 221 28 307 67 312 468 176 103 642 105537) Roundto 3 de inal places What is the population standard deviation? ơ . What is the mean? p.358 143 According to the Range Rule of Thumb descrbed n he videos ofchaptea, 95% of an onurtis would have bet on what ninal of re? ninimum D maximum 729 217 Seminole county covers 345 square miles Would Seminole county be consiered an outler if it was located in New Jersey? A. No. 345 is outside the interval of usual values B. Yes, 345 is within the interval of usual values C. Yes, 345 is outside the interval of usual values D. No, 345 is within the interval of usual values co 2 ick to select your answer(s)
Answer #1

1038 381585 2 2 2 2 7 )737333373373737773377 1 04 1 64 0 4 3 6 6 5 8 4 7 52 09 3 ツ 2 24 2 12345678901 10 12 13 14 15 16 17 18

from above mean mu =358.143

and std deviation sigma =185.537

minimum =358.143-2*185.537=-12.931 ~ 0 (as area can not be negative)

maximum=358.143+2*185.537=729.217

option D is correct; No ; 345 is with in interval of usual values.